Henry invested $12,000 in a new restaurant. When the restaurant sold two years later, he received $27,000. Find his average annual return by solving the equation 12,000(I +r)^2= 27,000.
Answer by checkley77(12844)

12,000(I +r)^2= 27,000.
12,000(1+r)^2=27,000
(1+r)^2=27,000/12,000
(1+r)^2=2.25
1+r=sqrt2.25
1+r=1.5
r=1.5-1
r=.5 or 50% is the annual increase in value.